# Viewpoint Matters: Dynamically Optimizing Viewpoints with Masked Autoencoder for Visual Manipulation

### v1

@@ Abstract
Robotic manipulation continues to be a challenge, and imitation learning (IL) enables robots to learn tasks from expert demonstrations. Current IL methods typically rely on fixed camera setups, where cameras are manually positioned in static locations, imposing significant limitations on adaptability and coverage. Inspired by human active perception, where humans dynamically adjust their viewpoint to capture the most relevant and least noisy information, we propose MAE-Select, a novel framework for active viewpoint selection in single-camera robotic systems. MAE-Select fully leverages pre-trained multi-view masked autoencoder representations and dynamically selects the next most informative viewpoint at each time chunk without requiring labeled viewpoints. Extensive experiments demonstrate that MAE-Select improves the capabilities of single-camera systems and, in some cases, even surpasses multi-camera setups. The project will be available at https://mae-select.github.io.
Index Terms—  Manipulation, Imitation Learning, Active Perception
@@ 1 Introduction
Robotic manipulation is a core challenge in robotics. Imitation Learning (IL) [17, 12, 6, 5, 2] has become a leading approach for enabling robots to learn complex tasks through expert demonstrations. Recently, advances in deep generative models [7, 10, 14, 9], have empowered IL by allowing robots to process high-dimensional sensory inputs, leading to promising results in robotic manipulation [3, 25, 1, 19].
However, most current IL methods rely on fixed camera setups, either single or multiple cameras, which pose significant limitations. In fixed single-camera setups [16, 18], though cost-effective, robots face challenges due to the limited field of view, which may obstruct critical parts of the environments, negatively impacting task performance. Multi-camera setups, while designed to provide more comprehensive scene coverage, introduce their own complexities: the abundance of redundant or irrelevant information can overwhelm learning algorithms and decrease efficiency. As shown in Sec.3, these passive static multi-view setups do not always provide the most task-relevant and cleanest information, leading to suboptimal decision-making.
@@
To this end, we introduce MAE-Select, a framework designed to actively select optimal viewpoints for single-camera robotic setups. MAE-Select first fully utilizes the powerful pre-trained representations from the multi-view masked autoencoders (MAEs) [8, 23], leveraging its complete encoder-decoder architecture to obtain estimated multi-view representations. Unlike prior works that focus on fixed viewpoints [21], MAE-Select dynamically predicts the next better viewpoint based on the current chunk of visual and action information. Crucially, this viewpoint selection is learned solely through imitation learning, which does not require manual labels for optimal views. It demonstrates significant potential in the advancement of single-camera robotic manipulation.
Fig. 1: Illustration of our proposed method. Left depicts the pre-training stage of the multi-view masked autoencoder with joint embeddings. Middle illustrates the training process of our framework using imitation learning. Right demonstrates how the framework operates during inference.
@@ 2 Method
We introduce MAE-Select, a novel framework that enables robotic agents to actively select informative viewpoints for manipulation tasks. The core of our approach is a new formulation in which optimal viewpoint selection is learned implicitly through an imitation learning objective, without requiring explicit supervision or reinforcement learning. Unlike previous work that only uses pre-trained encoders [20, 21], our method exploits both encoder and decoder of a multiview masked autoencoder (MV-MAE), enabling the agent to construct a rich scene representation from a single view, crucial for accurate action prediction and informed viewpoint selection. An overview of our architecture is shown in Fig.1.
@@ 2.1 Problem Formulation
In our active perception setting, the policy does not have access to all viewpoints simultaneously. Instead, it operates on a single selected viewpoint $o_{t}^{v_{t}}$ at time $t$. Therefore, the objective is to learn two key policies:
@@ 2.2 Multi-View Masked Autoencoder Pre-training
We pre-train a Multi-View Masked Autoencoder (MV-MAE) on demonstration data to learn a compact, 3D-aware scene representation (Fig. 1, left).
Given a multi-view observation $O_{t}$, we extract feature maps $\{\bar{o}_{t}^{v}\}_{v\in V}$ and apply dual masking: (1) patch masking, which randomly masks a large portion of patches within each view, and (2) view masking, which randomly drops entire views to encourage cross-view reasoning. The remaining patches, with view and position embeddings, are fed into a Transformer encoder $f_{\phi}$ to produce latent $z_{t}^{m}$. A decoder $g_{\phi}$ reconstructs all views from $z_{t}^{m}$, mask tokens, and the robot state embedding $s_{t}$ by minimizing
$\mathcal{L}_{\text{MAE}}=\mathbb{E}_{O_{t},s_{t}}\left[\|O_{t}-g_{\phi}(f_{\phi}(O_{t}^{m}),s_{t})\|^{2}_{2}\right].$ (1)
This pre-training enables the model to infer full 3D scenes from partial or occluded inputs for downstream tasks.
@@ 2.3 Next Better Viewpoint Selection
Following pre-training, we jointly fine-tune the model and train both the action and viewpoint selection policies (Fig. 1, middle). Although action decoder training uses ground-truth actions, defining a ground-truth “optimal” viewpoint is impractical, as it depends on task and context. Instead of reinforcement learning, which is difficult due to the sim-to-real gap, we propose a label-free strategy to train the viewpoint selector $\pi_{\psi}$. Specifically, the action prediction loss on a future data chunk $\mathcal{D}_{t+T}$ is used as a supervisory signal for the viewpoint chosen in the current chunk $\mathcal{D}_{t}$. We apply this process to two consecutive trajectory chunks, $\mathcal{D}_{t}$ and $\mathcal{D}_{t+T}$.
Processing the Current Chunk ($\mathcal{D}_{t}$): For the current chunk starting at time $t$, we begin with a single view $o_{t}^{v}$, selected uniformly at random. This single-view observation is passed through the entire pre-trained MV-MAE ($f_{\phi},g_{\phi}$) to generate an estimated multi-view feature context $C_{t}=g_{\phi}(f_{\phi}(o_{t}^{v}),s_{t})$. This context $C_{t}$ is then fed into the action decoder $\pi_{\theta}$, a diffusion-based policy, which is trained to predict the noise $\epsilon_{t}$ added to the expert action trajectory $a_{t:t+T-1}$. The action loss for this chunk is computed as:
@@
Selecting the View for the Next Chunk ($\mathcal{D}_{t+T}$): The viewpoint selector $\pi_{\psi}$, implemented as a Transformer encoder, takes the feature context $C_{t}$ and the ground-truth action trajectory $a_{t:t+T-1}$ from the current chunk to predict a probability distribution of viewpoints for the next chunk:
$\hat{c}=\text{Softmax}(\pi_{\psi}(C_{t},a_{t:t+T-1}))$ (3)
To select a viewpoint while maintaining differentiability for backpropagation, we employ a straight-through estimator as VQ-VAE [24]. In the forward, we perform an argmax operation to obtain one-hot vector $\mathbf{y}$ representing the chosen view. For the backward, gradients flow through the continuous softmax probabilities $\hat{c}$. This is implemented as:
[…]
@@
Processing the Next Chunk ($\mathcal{D}_{t+T}$): The one-hot vector $\mathbf{y}$ is used to select a single viewpoint $o_{t+T}^{\hat{v}}$ from the next observation set $O_{t+T}$. This selection is performed via a dot product $o_{t+T}^{\hat{v}}=\mathbf{y}\cdot O_{t+T}$, which also enables gradient flow back to the selection policy. This selected view is then processed through the identical pipeline as the first chunk, yielding a corresponding action loss, $\mathcal{L}_{\text{action}}^{(t+T)}$.
Updating the Viewpoint Selector: There is no explicit “view loss.” The viewpoint selector $\pi_{\psi}$ is trained using gradients from the future action loss $\mathcal{L}_{\text{action}}^{(t+T)}$, backpropagated through the action decoder and selected observation via STE, directly optimizing $\pi_{\psi}$ to select viewpoints that minimize future action prediction error.
The overall objective combines both action losses and the auxiliary reconstruction loss:
@@
where $\pi_{\theta}$ is updated by both action losses, while $\pi_{\psi}$ is optimized only by the future action loss.
Inference: During inference (Fig. 1, right), the process is autoregressive. Starting from a random initial viewpoint, it predicts both the current action chunk and the optimal viewpoint for the next chunk. The predicted action is then used alongside the newly selected viewpoint for the subsequent step, creating a dynamic perception-action loop.

### v2

@@ Abstract
Robotic manipulation continues to be a challenge, and imitation learning (IL) enables robots to learn tasks from expert demonstrations. Current IL methods typically rely on fixed camera setups, where cameras are manually positioned in static locations, imposing significant limitations on adaptability and coverage. Inspired by human active perception, where humans dynamically adjust their viewpoint to capture the most relevant and least noisy information, we propose MAE-Select, a novel framework for active viewpoint selection in single-camera robotic systems. MAE-Select fully leverages pre-trained multi-view masked autoencoder representations and dynamically selects the next most informative viewpoint at each time chunk without requiring labeled viewpoints. Extensive experiments demonstrate that MAE-Select improves the capabilities of single-camera systems and, in some cases, even surpasses multi-camera setups. The project will be available at https://sites.google.com/view/mae-select.
@@ 1 Introduction
Figure 1: Comparison of visual perception paradigms for robotic manipulation. Traditional passive static setups (left), such as fixed single-camera or multi-camera systems, often suffer from visual occlusions and data redundancy. In contrast, our proposed MAE-Select framework (right) employs active dynamic perception to efficiently acquire optimal task-relevant viewpoints.
Robotic manipulation is a core challenge in robotics, critical to applications ranging from industrial automation to healthcare. Imitation Learning (IL) [pomerleau1988alvinn, jang2022bc, florence2022implicit, ebert2021bridge, brohan2022rt, liu2024foam] has become a leading approach for enabling robots to learn complex tasks through expert demonstrations. Recently, advances in deep generative models [goodfellow2020generative, huang2023diffusion], such as variational autoencoders (VAEs) [kingma2013auto] and diffusion models [ho2020denoising], have empowered IL by allowing robots to process high-dimensional sensory inputs, such as images, leading to promising results in robotic manipulation [chi2023diffusion, zhao2023learning, bharadhwaj2024roboagent, reuss2023goal].
However, most current IL methods rely on fixed camera setups, either single or multiple cameras, which pose significant limitations. In fixed single-camera setups [pari2021surprising, qin2022one], though practical and cost-effective, robots face challenges due to the limited field of view, which may obstruct critical parts of the environment or objects, negatively impacting task performance. Multi-camera setups, while designed to provide more comprehensive scene coverage, introduce their own complexities: the abundance of redundant or irrelevant information can overwhelm learning algorithms and decrease efficiency. As shown in Sec.4, these passive static multi-view setups do not always provide the most task-relevant and cleanest information, leading to inefficiencies and suboptimal decision-making.
@@
To this end, we introduce MAE-Select, a framework designed to actively select optimal viewpoints for single-camera robotic setups. MAE-Select first fully utilizes the powerful pre-trained representations from the multi-view masked autoencoders (MAEs) [he2022masked, tong2022videomae], leveraging its complete encoder-decoder architecture to obtain estimated multi-view representations. Unlike prior works that focus on fixed viewpoints [seo2023multi], MAE-Select dynamically predicts the next better viewpoint based on the current chunk of visual and action information. Crucially, this viewpoint selection is learned solely through imitation learning, which does not require manual labels for optimal views. It demonstrates significant potential in the advancement of single-camera robotic manipulation.
@@
2 Related Work
2.1 Imitation Learning for Manipulation
Imitation learning (IL), which enables agents to acquire complex skills by observing and mimicking expert behavior, has become a cornerstone approach for robotic manipulation [avigal2022speedfolding, arunachalam2023holo, de2019causal, duan2017one, jang2022bc, johns2021coarse, wen2022catgrasp, li2025splatter, han2026fsag]. A classic method within this paradigm is behavioral cloning (BC) [pomerleau1988alvinn], which formulates the problem as supervised learning by directly mapping observations to actions. Despite its simplicity, BC remains widely used due to its ease of implementation and scalability.
Recent years have seen a surge of IL techniques leveraging high-capacity neural networks and large-scale visual observations, driven by advances in deep learning and data collection platforms. Several methods now incorporate high-dimensional visual modalities such as RGB images [florence2019self, rahmatizadeh2018vision, wen2022you, zhao2023learning, chi2023diffusion], 3D point clouds [ze20243d, chen2021unsupervised], and virtual reality demonstrations [zhang2018deep], allowing policies to learn rich perceptual representations. For instance, ACT [zhao2023learning] introduces a transformer-based action chunking strategy that improves sample efficiency and generalization. Meanwhile, Diffusion Policy [chi2023diffusion] formulates policy learning as a conditional denoising process, achieving stable and expressive visuomotor behavior across a range of complex tasks.
Despite these advances, most IL frameworks operate under fixed-camera configurations, limiting their adaptability to spatial variations and occlusions. While recent works have investigated multi-view learning or bird’s-eye views, these approaches often require multiple calibrated cameras or rely on extensive 3D reconstruction pipelines [shridhar2023perceiver]. In contrast, our work emphasizes the importance of selecting informative viewpoints from a single camera, seeking to bridge the gap between rich multi-view information and practical deployment constraints in manipulation tasks.
2.2 Active Perception and Next Best View Planning
The concept of active perception, famously pioneered by Bajcsy [bajcsy1988active], posits that an agent should dynamically alter its sensor state to improve perceptual understanding. In robotics, this is often formulated as the Next Best View (NBV) problem, where the goal is to determine the optimal sensor placement for tasks such as 3D reconstruction, object recognition, or navigation [connolly1985determination, delmerico2018comparison]. Traditional NBV approaches rely on information-theoretic metrics, such as minimizing entropy or maximizing expected information gain based on volumetric representations like octrees [potthast2014probabilistic].
Recently, with the rise of deep reinforcement learning (RL) and neural rendering, active vision has been applied to dynamic environments. Works utilizing Neural Radiance Fields (NeRF) [mildenhall2021nerf] or 3D Gaussian Splatting [kerbl20233d] have explored active camera trajectories to minimize rendering uncertainty [pan2022activenerf]. Furthermore, RL has been employed to learn camera control policies that maximize task-specific rewards [zhu2025active]. However, these methods often require complex reward engineering or extensive 3D reconstruction pipelines. In contrast, MAE-Select implicitly learns the optimal viewpoint selection through an imitation learning objective, directly tying active perception to the downstream manipulation task without requiring explicit NBV heuristics or ground-truth viewpoint labels.
2.3 Unsupervised Representation Learning
Unsupervised representation learning aims to extract semantically meaningful features from unlabeled data, enabling downstream tasks to benefit from improved generalization and data efficiency [sohn2015learning, jing2020self, donahue2016adversarial]. Among the leading approaches, Masked Autoencoders (MAE) [he2022masked] have emerged as a powerful technique, training models to reconstruct masked portions of the input by learning context-aware latent representations. This pretext task has led to significant advances in vision applications, including image classification, segmentation, and video understanding [li2022uniform, zhang2022hivit, tong2022videomae].
In the context of robotics, several works have incorporated MAE-based frameworks to improve policy learning from visual inputs. R3M [nair2022r3m] and Real-world MAE [radosavovic2023real] show that unsupervised pretraining can provide robust and generalizable representations for manipulation. Specifically, MV-MWM [seo2023multi] extends MAE to multi-view settings by encoding multiple perspectives during training and applying single-view inference at test time, demonstrating enhanced performance on several manipulation benchmarks. However, its viewpoint remains static, using predefined camera without exploring alternative viewpoints. In contrast, our work explicitly investigates the viewpoint selection problem, aiming to discover the most informative camera view for manipulation tasks.
Figure 2: Illustration of our proposed method. Left depicts the pre-training stage of the multi-view masked autoencoder with joint embeddings. Right illustrates the training process of our framework using imitation learning.
@@ 3 Method
We introduce MAE-Select, a novel framework that enables robotic agents to actively select informative viewpoints for complex manipulation tasks. The core of our approach is a new formulation in which optimal viewpoint selection is learned implicitly through an imitation learning objective—without requiring explicit supervision or reinforcement learning. In contrast to prior work that relies solely on the encoder of pre-trained models [seo2023masked, seo2023multi], our method leverages the full generative capacity of a multi-view masked autoencoder (MV-MAE), including both its encoder and decoder. This enables the agent to build a rich, 3D-aware representation of the scene from just a single viewpoint, which is essential for accurate action prediction and informed viewpoint selection. An overview of our architecture is shown in Fig.2. We begin by formalizing the problem, then describe our representation learning strategy, and finally explain how the action and viewpoint selection policies are trained jointly.
@@ 3.1 Problem Formulation
In our active perception setting, the policy does not have access to all viewpoints simultaneously. Instead, it operates on a single selected viewpoint $o_{t}^{v_{t}}$ at time $t$. Therefore, the objective is to learn a composite policy $\pi$ that can be decomposed into two key components:
@@ 3.2 Multi-View Masked Autoencoder Pre-training
To build a strong foundation for visual understanding, we pre-train a Multi-View Masked Autoencoder (MV-MAE) on the demonstration data. This pre-training stage, depicted in the left panel of the pipeline diagram, is designed to learn a compressed yet comprehensive representation of the multi-view scene.
The architecture is trained to reconstruct a full multi-view image set from a heavily masked input. Given a multi-view observation $O_{t}$, we first extract convolutional feature maps $\{\bar{o}_{t}^{v}\in\mathbb{R}^{H\times W\times D_{feat}}\}_{v\in V}$. We then apply a dual-masking strategy:
  • Patch Masking: A large percentage of the feature patches within each view’s map are randomly masked (i.e., zeroed out).
  • View Masking: Entire views are randomly masked, forcing the model to infer inter-view relationships.
The set of unmasked feature patches, augmented with learned 1D view-specific embeddings and fixed 2D sinusoidal position embeddings, are fed into a Transformer-based encoder, $f_{\phi}$, to produce a latent representation $z_{t}^{m}$. The decoder, $g_{\phi}$, then takes this latent code, a full set of mask tokens, and the robot’s joint state embedding $s_{t}$ to reconstruct the original, unmasked features for all viewpoints. The pre-training objective is to minimize the reconstruction error:
$\mathcal{L}_{\text{MAE}}(\phi)=\mathbb{E}_{O_{t},s_{t}\sim\mathcal{D}}\left[\|O_{t}-g_{\phi}(f_{\phi}(O_{t}^{m}),s_{t})\|^{2}_{2}\right]$ (1)
where $O_{t}^{m}$ denotes the masked input views. This process endows the model with a powerful generative prior, enabling it to hallucinate a full 3D scene representation from a single, potentially occluded input viewpoint during the downstream task.
@@ 3.3 Next Better Viewpoint Selection
Following pre-training, we fine-tune the model and train the action and viewpoint selection policies concurrently, as illustrated in the middle panel of the diagram. Our central contribution is a training strategy for the viewpoint selector $\pi_{\psi}$ without explicit labels for the "best" viewpoint. We achieve this by using the action prediction loss of a future time chunk as a supervisory signal for the viewpoint choice made in the current chunk. This process unfolds over two consecutive chunks of data from a trajectory, $\mathcal{D}_{t}$ and $\mathcal{D}_{t+T}$.
Algorithm 1 Training MAE-Select (Single Iteration)
0: Expert demonstration dataset $\mathcal{D}$, Pre-trained MV-MAE encoder $f_{\phi}$ and decoder $g_{\phi}$, Action policy $\pi_{\theta}$, Viewpoint selector $\pi_{\psi}$, Chunk size $T$.
1: Sample consecutive data chunks $(\mathcal{D}_{t},\mathcal{D}_{t+T})$ from $\mathcal{D}$.
2:// Step 1: Process Current Chunk
3: Sample an initial viewpoint $v_{t}\sim\mathcal{U}(V)$.
4: Extract single-view observation $o_{t}^{v_{t}}$ from $\mathcal{D}_{t}$.
5: Compute multi-view context: $C_{t}=g_{\phi}(f_{\phi}(o_{t}^{v_{t}}),s_{t})$.
6: Sample noise $\epsilon_{t}$ and timestep $k\sim\mathcal{U}(1,K)$.
7: Compute action loss: $\mathcal{L}_{\text{action}}^{(t)}=\|\epsilon_{t}-\pi_{\theta}(C_{t},a_{t:t+T-1}^{0}+\epsilon_{t},k)\|^{2}$.
8:// Step 2: Select Next Viewpoint via STE
9: Predict viewpoint probabilities: $c_{\text{hat}}=\text{Softmax}(\pi_{\psi}(C_{t},a_{t:t+T-1}))$.
10: Discretize to one-hot vector: $\mathbf{y}_{\text{hard}}=\text{argmax}(c_{\text{hat}})$.
11: Apply Straight-Through Estimator: $\mathbf{y}=c_{\text{hat}}+\text{sg}[\mathbf{y}_{\text{hard}}-c_{\text{hat}}]$.
12:// Step 3: Process Next Chunk
13: Select next observation using $\mathbf{y}$: $o_{t+T}^{\hat{v}}=\mathbf{y}\cdot O_{t+T}$.
14: Compute new context: $C_{t+T}=g_{\phi}(f_{\phi}(o_{t+T}^{\hat{v}}),s_{t+T})$.
15: Sample new noise $\epsilon_{t+T}$ and timestep $k^{\prime}$.
16: Compute next action loss: $\mathcal{L}_{\text{action}}^{(t+T)}=\|\epsilon_{t+T}-\pi_{\theta}(C_{t+T},a_{t+T:t+2T-1}^{0}+\epsilon_{t+T},k^{\prime})\|^{2}$.
17:// Step 4: Optimization
18: Compute total loss: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{action}}^{(t)}+\lambda_{1}\mathcal{L}_{\text{action}}^{(t+T)}+\lambda_{2}\mathcal{L}_{\text{MAE}}$.
19: Update $\theta,\phi$ using $\nabla\mathcal{L}_{\text{total}}$.
20: Update $\psi$ using gradients flowing back from $\mathcal{L}_{\text{action}}^{(t+T)}$ via STE.
Processing the Current Chunk ($\mathcal{D}_{t}$): For the current chunk starting at time $t$, we begin with a single view $o_{t}^{v}$, selected uniformly at random. This single-view observation is passed through the entire pre-trained MV-MAE ($f_{\phi},g_{\phi}$) to generate an estimated multi-view feature context $C_{t}=g_{\phi}(f_{\phi}(o_{t}^{v}),s_{t})$. This context $C_{t}$ is then fed into the action decoder $\pi_{\theta}$—a diffusion-based policy—which is trained to predict the noise $\epsilon_{t}$ added to the expert action trajectory $a_{t:t+T-1}$. The action loss for this chunk is computed as:
@@
Selecting the View for the Next Chunk ($\mathcal{D}_{t+T}$): The viewpoint selector $\pi_{\psi}$, implemented as a Transformer encoder, takes the feature context $C_{t}$ and the ground-truth action trajectory $a_{t:t+T-1}$ from the current chunk to predict a probability distribution over the available viewpoints for the next chunk:
$c_{\text{hat}}=\text{Softmax}(\pi_{\psi}(C_{t},a_{t:t+T-1}))$ (3)
To select a discrete viewpoint while maintaining differentiability for backpropagation, we employ a straight-through estimator (STE) like VQ-VAE [van2017neural]. In the forward pass, we perform an argmax operation to get a one-hot vector $\mathbf{y}_{\text{hard}}$ representing the chosen view. For the backward pass, gradients flow through the continuous softmax probabilities $c_{\text{hat}}$. This is implemented as:
$\mathbf{y}=c_{\text{hat}}+sg[\mathbf{y}_{\text{hard}}-c_{\text{hat}}]$ (4)
@@
Processing the Next Chunk ($\mathcal{D}_{t+T}$): The one-hot vector $\mathbf{y}$ is used to select a single viewpoint $o_{t+T}^{\hat{v}}$ from the next observation set $O_{t+T}$. This selection is performed via a dot product $o_{t+T}^{\hat{v}}=\mathbf{y}\cdot O_{t+T}$, which also enables gradient flow back to the selection policy. This selected view is then processed through the identical pipeline as the first chunk, yielding a new feature context $C_{t+T}$ and a corresponding action loss, $\mathcal{L}_{\text{action}}^{(t+T)}$.
Updating the Viewpoint Selector: There is no explicit "View Loss" term. Instead, the training signal for the viewpoint selector $\pi_{\psi}$ is implicitly provided by the action loss of the subsequent time chunk, $\mathcal{L}_{\text{action}}^{(t+T)}$. Through the use of the STE, gradients from $\mathcal{L}_{\text{action}}^{(t+T)}$ are backpropagated through the action decoder and the selected observation to update the parameters $\psi$ of the selection model. This process directly optimizes the selector to choose viewpoints that minimize future action prediction error, a mechanism visually represented by the "View Loss" feedback loop in the training diagram.
The comprehensive training objective, therefore, combines the losses from both time chunks with the auxiliary reconstruction objective to train all components end-to-end:
@@
By optimizing this single objective, the action policy $\pi_{\theta}$ is updated by both action loss terms, while the viewpoint selector $\pi_{\psi}$ is specifically optimized by the gradients flowing from the future action loss, $\mathcal{L}_{\text{action}}^{(t+T)}$.
Figure 3: Overview of the inference pipeline for the proposed framework.
Inference: During inference, as shown in Fig.3, the process is autoregressive. Starting with a random initial viewpoint, the agent uses the observation to predict both the first action chunk and the optimal viewpoint for the next chunk. The predicted action is then used alongside the newly selected viewpoint for the subsequent step, creating a dynamic perception-action loop.
Figure 4: The viewpoint settings for various robotic tasks, showcasing the viewpoints used to evaluate performance across different simulation and real-world scenarios.
4 Experiments
We evaluate MAE-Select across 3 challenging scenarios and 8 demanding tasks in simulation, including simulations in ACT [zhao2023learning], RLBench [james2019rlbench], and our designed robot in MuJoCo [todorov2012mujoco], along with three real-world tasks.
4.1 Tasks
The 11 tasks cover various scenarios, providing a comprehensive evaluation of our method. The viewpoint settings are shown in Fig. 4. We categorize these tasks across four different environments:
4.1.1 ACT Simulations [zhao2023learning]:
We focus on the Bimanual Insertion task, where the robot’s arms pick up a socket and a peg for a mid-air insertion.
4.1.2 RLBench [james2019rlbench]:
We select four varied tasks as follows:
  • Phone On Base: The robot needs to pick up and place a phone onto its base.
  • Pick Up Cup: The robot needs to grasp and lift a specific cup from among several cups.
  • Unplug Charger: The robot is tasked with removing a charger from a socket, involving careful manipulation to avoid damaging delicate components.
  • Take Umbrella Out Of Stand: This task requires the robot to pick up a small umbrella.
4.1.3 MuJoCo [todorov2012mujoco]:
We design three customized tasks for our robot:
  • Put Box In Cabinet: The robot needs to pick up a box and place it inside a cabinet, requiring precise spatial reasoning.
  • Put Box In Bin: In this task, the robot places a box into a bin, testing its ability to interact with constrained spaces.
  • Put Box In Bin with Disturbance: Similar to the previous task, but the robot must select a specific block from multiple blocks.
4.1.4 Real-World Evaluation:
We designed three tasks similar to the simulation setups but involving real objects: Put Bitter Melon In Cabinet, Put Eggplant To Bowl, and Put Eggplant To Bowl with Disturbance.
4.2 Data Collection
Following previous methods [zhao2023learning, james2019rlbench], we collect demonstrations with a scripted policy for each task in simulation. The datasets capture the robot’s proprioceptive data, action sequences, and multi-view visual observations. For real-world data acquisition, we implement a low-cost teleoperation platform inspired by Gello [wu2024gello]. Our robot consists of a 7DOF Ufactory xarm 7 robotic arm and a parallel-jaw gripper. Our camera setup includes two statically mounted (top and left) and one wrist-mounted Realsense D435 camera. For each task, we collected 50 expert demonstrations, while the initial placement of the objects is varied uniformly in 2D regions. The final dataset comprises RGB data from two cameras with a resolution of 224×224, along with joint states and action sequences.
4.3 Implementation
We explore a single-camera control setup where the system is trained with multiple camera views but deployed with a single moving camera. At each time chunk, the camera is placed at one of the training viewpoints, enabling flexible operation in real-world settings where only one camera is available. We base our implementation on the architecture of the diffusion policy [chi2023diffusion]. The action space corresponds to the joint angles of the robot arm, while the image observations have a resolution of $224\times 224$ with a patch size of 16. Our masked autoencoder utilizes a 12-layer ViT [dosovitskiy2020image] for encoder and an 8-layer ViT for decoder, with an embedding dimension of 512. During pretraining, we use a batch size of 128 over 100 epochs. For RLBench and real-world tasks, the time chunk size is set to 20, with a total of 600 epochs. For other tasks, the time chunk size is 100, with 1,000 epochs in total. $\lambda_{1}$ and $\lambda_{2}$ are empirically set to 2.0 and 1.0, respectively. The batch size is set to 32, and the training undergoes two stages, where only the imitation component is trained during the first half of the epochs, and both imitation and view selection components are trained in the second half. For each method, we evaluate the best-performing checkpoints from the last three evaluated at 100-intervals, with 50 environment initializations in simulation. All models were trained and tested on NVIDIA RTX 4090 GPUs.